OpenStudy (swissgirl):

Let \(f(A,*) \implies (B, \cdot)\) be homomorphic. Prove that if \(f^{-1}\)is a function then it is homomorphic

OpenStudy (anonymous):

work directly from the definition the hard part is probably figuring out what you need to show, showing it is straight forward do you know what you need to do?

OpenStudy (swissgirl):

Well i need to show that it is a bijection?

OpenStudy (anonymous):

oh no you only need to show what it asks you to show that the inverse function is a homomorphism you do not need to show that it is a bijection and in fact it need not be because you only know \(f\) is one to one, not necessarily onto

OpenStudy (swissgirl):

well it must be onto in order for the inverse to be a function

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

it just has to be one to one

OpenStudy (swissgirl):

ohhhhh noooooooooooooooo

OpenStudy (swissgirl):

idk i read that somewhere in my bookkk guess i am wrong

OpenStudy (anonymous):



OpenStudy (anonymous):

think of the map that send \(n\mathbb{Z}\to \mathbb{Z}\) via \(a\to a\)

OpenStudy (swissgirl):

wait if its not onto then only the range of F can be mapped to A

OpenStudy (anonymous):

that is one to one, but certainly not onto you are only being asked to show that IF \(f\) is a homo and IF \(f\) is one to one (so the inverse exists) THEN \(f^{-1}\) is a homo

OpenStudy (swissgirl):

righhhttttttt

OpenStudy (anonymous):

that is right only the range can be mapped to A

OpenStudy (anonymous):

so the domain of \(f^{-1}\) is the range of \(f\), not necessarily all of \(B\)

OpenStudy (swissgirl):

ohhhhhhh okkkkkkkkkkkkkk

OpenStudy (anonymous):

come up with some examples for yourself here is one, map \(\mathbb{R}\) to \(\mathbb{R}\times \mathbb{R}\) via \(x\to (x,0)\) this is certainly one to one, and invertible, but it is not onto for sure

OpenStudy (swissgirl):

yaaa

OpenStudy (anonymous):

in any case your job is to show that \(f^{-1}\) is a homo, that is all you get to assume it exists, so you get to assume that \(f\) is one to one, and also you are given that \(f\) is homo

OpenStudy (anonymous):

write down exactly what you have to show, and exactly what you know

OpenStudy (swissgirl):

right so that means that \( f(x_1*x_2)=f(x_1)*f(x_2) \)

OpenStudy (anonymous):

careful

OpenStudy (swissgirl):

gotta give me a sec to see how i can pull it all together

OpenStudy (swissgirl):

ya there shld be that bullet

OpenStudy (swissgirl):

ok ill rewrite it

OpenStudy (anonymous):

\(*\) is the operation in A, \(\cdot\) is the operation in B

OpenStudy (swissgirl):

\( f(x_1*x_2)=f(x_1) \cdot f(x_2) \)

OpenStudy (anonymous):

you get to assume that, yes you also get to assume one more thing about \(f\)

OpenStudy (swissgirl):

now \(f^{-1}(y_1*y_2)=f^{-1}(y_1) \cdot f^{-1}(y_2) \)

OpenStudy (swissgirl):

can we assume that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

be careful about where things live \(x\in A,f(x)\in B\)

OpenStudy (swissgirl):

\(f^{-1}(y_1)=x_1 \)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i assume you mean "let \(f^{-1}(y_1)=x_1\)"

OpenStudy (swissgirl):

yaaaaaaaaaaaa

OpenStudy (swissgirl):

Let \( f^{-1}(y_1)=x_1 \) and\( f^{-1}(y_2)=x_2 \)

OpenStudy (anonymous):

don't forget we know one more thing about \(f\) namely that IF \(f(x)=f(x_2)\) THEN \(x_1=x_2\) we may need that later

OpenStudy (anonymous):

ok good now what exactly do you want to show about this

OpenStudy (swissgirl):

hahah ok this may be incorrect \( f(f^{-1}(y_1)*f(^{-1}(y_2)) = f(f^{-1}(y_1)) \cdot f(f^{-1}(y_2)) \)

OpenStudy (swissgirl):

is this totally off track?

OpenStudy (anonymous):

i think i is over complicated write down exactly what it means for \(f^{-1}\) to be a homo, being careful with the multiplication symbols then i think we can show it, once we know exactly what we need to show

OpenStudy (anonymous):

in other words, as usual with a proof, write down precisely what is needed to be shown, working directly from the definition in other words, translate "\(f^{-1}\) is homomorphic" in to math

OpenStudy (swissgirl):

ok we need to show that \(f^{-1}(y_1 \cdot y_2) = f^{-1}(y_1)*f^{-1}(y_2) \)

OpenStudy (experimentx):